# Relativistic quintuple-zeta basis sets for the s block

### v1

@@ Abstract
Relativistic basis sets of quintuple-zeta quality are presented for the s-block elements. The basis sets include SCF exponents for the occupied spinors and for the np shell (the latter is considered here a valence orbital). Valence and core correlating functions were optimized within multireference SDCI calculations for the ground valence configuration. Diffuse functions optimized for the corresponding anions or derived from neighboring elements are also provided. The new basis sets were applied to a range of basic atomic and molecular properties for benchmarking purposes. Smooth convergence to the basis set limit is observed with increased basis set quality from existing double-zeta, triple-zeta, and quadruple-zeta to the newly developed quintuple-zeta basis sets. Use of these basis sets in combination with state-of-the-art approaches for treatment of relativity and correlation will allow achieving higher accuracy and lower uncertainty than previously possible in calculations on heavy atoms and molecules. The basis sets are available from zenodo.org.
Gaussian basis sets, relativistic basis sets, p elements, quintuple zeta, correlating functions
@@ I Introduction
For heavy-element calculations, the Dyall basis sets Dyall (1998, 2002, 2004, 2006, 2007a, 2007b, 2009); Dyall and Gomes (2010); Gomes, Dyall, and Visscher (2010); Dyall (2011, 2012a, 2012b, 2016) are commonly used, as they cover all elements up to Z=118 at the double-zeta (2z), triple-zeta (3z), and quadruple-zeta (4z) levels. The correlation energy CBS limit can currently be obtained from extrapolation of the 3z and 4z results (as 2z basis sets are insufficient for an accurate description of correlation effects). The extrapolation commonly uses one of a few schemes, such as those presented in Refs. Helgaker et al., 1997; Lesiuk and Jeziorski, 2019; Martin, 1996a; Amir Karton, 2006. However, this extrapolation is often shown to be the limiting factor in the accuracy of the calculated energies and properties Leimbach et al. (2020); Skripnikov (2021); Haase et al. (2020).
In order to overcome this limitation and increase the accuracy even further, basis sets of quintuple-zeta (5z) quality are needed. This paper reports the development and testing of 5z quality basis sets for the s-block elements. This work continues the ongoing effort Dyall (1998, 2002, 2004, 2006, 2007a, 2007b, 2009); Dyall and Gomes (2010); Gomes, Dyall, and Visscher (2010); Dyall (2011, 2012a, 2012b, 2016) to provide high-quality basis sets for all-electron correlated relativistic calculations across the periodic table and in particular for the heaviest elements. In addition to the SCF sets, which include functions to describe the ns and np shells, the basis sets include high-angular momentum correlating functions for valence and core shells, functions for dipole polarization for the outer core shells, and diffuse functions. The performance of these basis sets is illustrated in calculations of basic atomic and molecular properties of the s-block elements and selected diatomic molecules.
@@ III Basis set results
A summary of the SCF basis sets is given in Table 1. These do not include functions in the occupied symmetries that are added for valence flexibility or correlation. The total SCF energies for these basis sets are given as a reference for validation. These energies were calculated as described above, for the specified isotope of each element, with the basis set extension of GRASP Dyall and Fægri (1996).
Table 1: SCF basis set sizes for the s-block elements, and total energies in hartrees for the ground state of each element with the shown isotope number.
Element Basis SCF energy Element Basis SCF energy00
7Li 24s14p 7.43353312-7.43353312 85Rb 40s33p18d 2979.80501372-2979.80501372
9Be 23s12p 14.57588777-14.57588777 88Sr 40s33p18d 3178.07996988-3178.07996988
23Na 28s18p 162.07808696-162.07808696 133Cs 43s38p23d 7786.77168269-7786.77168269
24Mg 28s18p 199.93506634-199.93506634 138Ba 43s38p23d 8135.64502828-8135.64502828
39K 35s26p 601.52595321-601.52595321 223Fr 43s43p25d16f 24308.19431157-24308.19431157
40Ca 35s26p 679.71016058-679.71016058 226Ra 43s43p25d16f 25028.18889163-25028.18889163
@@ IV Applications
To investigate the performance of the newly-developed basis sets, calculations were carried out on several atomic and molecular systems. In order to show the consistency of the new basis sets with the previous ones, we compared calculations with the 5z basis sets to those obtained with the 2z, 3z and 4z basis sets. Additionally, extrapolation to the complete basis set (CBS) limit from different quality basis sets is demonstrated.
@@
The CBS(4z) limit was determined from calculations using up to the quadruple-zeta basis set, while CBS(5z) also includes the quintuple-zeta basis set results in the extrapolation procedure. The SCF energies are not extrapolated, because they are converged sufficiently for the larger basis sets, i.e. $E^{\text{SCF}}_{\text{CBS(4z)}}$ = $E^{\text{SCF}}_{\text{4z}}$ and $E^{\text{SCF}}_{\text{CBS(5z)}}$ = $E^{\text{SCF}}_{\text{5z}}$. The correlation energies are extrapolated using the scheme proposed by Martin Martin (1996b),
@@ IV.1 Ionization potentials and electron affinities
H CCSD 0.7010 0.7376 0.7467 0.7509 0.7544 0.7547
0.75420375(3) Andersen, Haugen, and Hotop (1999)
@@
CCSD(T) 0.5287 0.5437 0.5464 0.5471 0.5483 0.5480
0.547926(25) Hotop and Lineberger (1985)
@@
CCSD(T) 0.4434 0.4764 0.4818 0.4841 0.4858 0.4861
0.485916(21) Frey, Breyer, and Hotop (1978)
@@
CCSD(T) 0.4269 0.4606 0.4674 0.4696 0.4723 0.4722
0.4715983(38) Navarro Navarrete et al. (2024)
@@
All the CBS(5z) IPs and EAs are in remarkably good agreement (a few meV) with experimental values. We do not expect perfect agreement with experiment, as core-correlation, excitations beyond perturbative triples, and higher order relativistic effects are missing from these calculations. Thus, the new basis sets provide an excellent quality basis for higher order corrections. For Fr no experimental EA is available, but our result is in excellent agreement with the relativistic Fock space coupled cluster value (0.491 eV) from Ref. Landau et al., 2001.
@@ IV.2 Bond lengths and dissociation energies
CCSD(T) 3.0907 3.0818 3.0788 3.0780 3.0763 3.0767
3.0788(7) Rosner, Holt, and Gaily (1975)
@@
CCSD(T) 3.9858 3.9335 3.9257 3.9228 3.9202 3.9200
3.9051 Huber and Herzberg
@@
CCSD(T) 4.2651 4.2230 4.2167 4.2139 4.2126 4.2119
4.2099 Amiot, Crozet, and Verges (1985)
@@
CCSD(T) 0.4001 0.5119 0.5338 0.5409 0.5490 0.5488
0.563(4) Vinogradov and Volyak (1967)
@@
CCSD(T) 1.7569 1.7478 1.7493 1.7505 1.7498 1.7507
1.7500 Huber and Herzberg
@@
CCSD(T) 2.0236 1.9800 1.9590 1.9574 1.9534 1.9544
1.967 Huber and Herzberg
@@
CCSD(T) 2.1299 2.0861 2.0804 2.0787 2.0773 2.0772
2.07537 Huber and Herzberg
@@
CCSD(T) 2.2633 2.1945 2.1823 2.1797 2.1784 2.1777
2.162 Huber and Herzberg
@@ V Conclusions
The basis sets will be made available on zenodo.org when this manuscript is published.

### v2

@@ Abstract
Relativistic basis sets of quintuple-zeta quality are presented for the s-block elements. The basis sets include SCF exponents for the occupied spinors and for the np shell (the latter is considered here a valence orbital). Valence and core correlating functions were optimized within multireference SDCI calculations for the ground valence configuration. Diffuse functions optimized for the corresponding anions or derived from neighboring elements are also provided. The new basis sets were applied to a range of basic atomic and molecular properties for benchmarking purposes. Smooth convergence to the basis set limit is observed with increased basis set quality from existing double-zeta, triple-zeta, and quadruple-zeta to the newly developed quintuple-zeta basis sets. Use of these basis sets in combination with state-of-the-art approaches for treatment of relativity and correlation will allow achieving higher accuracy and lower uncertainty than previously possible in calculations on heavy atoms and molecules. The basis sets are available at https://doi.org/10.5281/zenodo.17088050.
@@ I Introduction
For heavy-element calculations, the Dyall basis sets Dyall (1998, 2002, 2004, 2006, 2007a, 2007b, 2009); Dyall and Gomes (2010); Gomes, Dyall, and Visscher (2010); Dyall (2011, 2012a, 2012b, 2016) are commonly used, as they cover all elements up to Z=118 at the double-zeta (2z), triple-zeta (3z), and quadruple-zeta (4z) levels. The correlation energy CBS limit can currently be obtained from extrapolation of the 3z and 4z results (as 2z basis sets are insufficient for an accurate description of correlation effects). The extrapolation commonly uses one of a few schemes, such as those presented in Refs. Helgaker et al., 1997; Lesiuk and Jeziorski, 2019; Martin, 1996; Amir Karton, 2006. However, this extrapolation is often shown to be the limiting factor in the accuracy of the calculated energies and properties Leimbach et al. (2020); Skripnikov (2021); Haase et al. (2020).
In order to overcome this limitation and increase the accuracy even further, basis sets of quintuple-zeta (5z) quality are needed. This paper reports the development and testing of 5z quality basis sets for the s-block elements. This work continues the ongoing effort Dyall (1998, 2002, 2004, 2006, 2007a, 2007b, 2009); Dyall and Gomes (2010); Gomes, Dyall, and Visscher (2010); Dyall (2011, 2012a, 2012b, 2016) to provide high-quality basis sets for all-electron correlated relativistic calculations across the periodic table and in particular for the heaviest elements. In addition to the SCF sets, which include functions to describe the ns and np shells, the basis sets include high-angular momentum correlating functions for valence and core shells, functions for dipole polarization for the outer core shells, and diffuse functions. The performance of these basis sets is illustrated in calculations of basic atomic and molecular properties of the s-block elements and selected diatomic molecules. Two of these molecules, BaF and RaF, are currently being investigated experimentally in the context of tests of the Standard Model, due to their relatively simple electronic structure and enhanced sensitivity to a variety of new physics phenomena Isaev, Hoekstra, and Berger (2010); Beloy et al. (2011); Safronova et al. (2018); eEDM collaboration et al. (2018); Altuntaş et al. (2018); Vutha, Horbatsch, and Hessels (2018); Kogel et al. (2025); Garcia Ruiz et al. (2020); Athanasakis-Kaklamanakis et al. (2025).
@@ III Basis set results
A summary of the SCF basis sets is given in Table 1. These do not include functions in the occupied symmetries that are added for valence flexibility or correlation. The total SCF energies for these basis sets are given as a reference for validation. These energies were calculated as described above, using the Gaussian nuclear charge distribution for the specified isotope of each element, with the basis set extension of GRASP Dyall and Fægri (1996).
Table 1: SCF basis set sizes for the s-block elements, and total energies in hartrees for the ground state of each element with the shown isotope number, compared to numerical energies.
Element Basis SCF (5z) SCF (numerical) Element Basis SCF (5z) SCF (numerical)
7Li 24s14p 7.4335331-7.4335331 7.4335331-7.4335331 85Rb 40s33p18d 2979.80501-2979.8050144 2979.805013-2979.805013
9Be 23s12p 14.57588-14.5758888 14.575892-14.575892 88Sr 40s33p18d 3178.0799-3178.07997070 3178.079969-3178.079969
23Na 28s18p 162.07808-162.0780877 162.078088-162.078088 133Cs 43s38p23d 7786.77168-7786.7716833 7786.771668-7786.771668
24Mg 28s18p 199.935066-199.935066 199.935067-199.935067 138Ba 43s38p23d 8135.645028-8135.645028 8135.645011-8135.645011
39K 35s26p 601.525953-601.525953 601.525954-601.525954 223Fr 43s43p25d16f 24308.19431-24308.1943122 24308.193350-24308.193350
40Ca 35s26p 679.71016-679.7101611 679.710161-679.710161 226Ra 43s43p25d16f 25028.18889-25028.1888922 25028.187810-25028.187810
@@ IV Applications
To investigate the performance of the newly-developed basis sets, calculations were carried out on several atomic and molecular systems. In order to show the consistency of the new basis sets with the previous ones, we compared calculations with the 5z basis sets to those obtained with the 2z, 3z and 4z basis sets. Additionally, extrapolation to the complete basis set (CBS) limit from different quality basis sets is demonstrated. We also compare the extrapolated results to experimental data where available, for reference purposes. It should be noted, however, that these calculations do not include the higher-order effects, such as excitations beyond perturbative triples and the Breit and QED corrections. The latter, in particular, are important for a meaningful comparison with experiment for heavy elements Leimbach et al. (2020); Guo et al. (2022). It should also be noted that the basis set energy can go below the numerical energy, due to the fact that the numerical DHF energy is not a rigorous lower bound Fægri, Jr. (2001).
@@
The CBS(4z) limit was determined from calculations using up to the quadruple-zeta basis set, while CBS(5z) also includes the quintuple-zeta basis set results in the extrapolation procedure. The SCF energies are not extrapolated, because they are converged sufficiently for the larger basis sets, i.e. $E^{\text{SCF}}_{\text{CBS(4z)}}$ = $E^{\text{SCF}}_{\text{4z}}$ and $E^{\text{SCF}}_{\text{CBS(5z)}}$ = $E^{\text{SCF}}_{\text{5z}}$. The correlation energies are extrapolated using the scheme proposed by Martin Martin (1996),
@@ IV.1 Ionization potentials and electron affinities
H CCSD 0.7010 0.7376 0.7467 0.7509 0.7544 0.7547 0.75420375(3) Andersen, Haugen, and Hotop (1999)
@@
CCSD(T) 0.5287 0.5437 0.5464 0.5471 0.5483 0.5480 0.547926(25) Hotop and Lineberger (1985)
@@
CCSD(T) 0.4434 0.4764 0.4818 0.4841 0.4858 0.4861 0.485916(21) Frey, Breyer, and Hotop (1978)
@@
CCSD(T) 0.4269 0.4606 0.4674 0.4696 0.4723 0.4722 0.4715983(38) Navarro Navarrete et al. (2024)
@@
All the CBS(5z) IPs and EAs are in remarkably good agreement (a few meV) with experimental values. We do not expect perfect agreement with experiment, as core-correlation, excitations beyond perturbative triples, and higher order relativistic effects are missing from these calculations. In particular, excitations beyond (T) have been shown to be very important for capturing the accurate values of electron affinities Pašteka et al. (2017); Leimbach et al. (2020); Guo et al. (2021). Thus, the new basis sets provide an excellent quality basis for higher order corrections. For Fr no experimental EA is available, but our result is in excellent agreement with the relativistic Fock space coupled cluster value (0.491 eV) from Ref. Landau et al., 2001.
@@ IV.2 Bond lengths and dissociation energies
CCSD(T) 3.0907 3.0818 3.0788 3.0780 3.0763 3.0767 3.0788(7) Rosner, Holt, and Gaily (1975)
@@
CCSD(T) 3.9858 3.9335 3.9257 3.9228 3.9202 3.9200 3.9051 Huber and Herzberg
@@
CCSD(T) 4.2651 4.2230 4.2167 4.2139 4.2126 4.2119 4.2099 Amiot, Crozet, and Verges (1985)
@@
CCSD(T) 0.4001 0.5119 0.5338 0.5409 0.5490 0.5488 0.563(4) Vinogradov and Volyak (1967)
@@
CCSD(T) 1.7569 1.7478 1.7493 1.7505 1.7498 1.7507 1.7500 Huber and Herzberg
@@
CCSD(T) 2.0236 1.9800 1.9590 1.9574 1.9534 1.9544 1.967 Huber and Herzberg
@@
CCSD(T) 2.1299 2.0861 2.0804 2.0787 2.0773 2.0772 2.07537 Huber and Herzberg
@@
CCSD(T) 2.2633 2.1945 2.1823 2.1797 2.1784 2.1777 2.162 Huber and Herzberg
@@ V Conclusions
The basis sets are available at https://doi.org/10.5281/zenodo.17088050.